OpenStudy (anonymous):

What is the slope and y-intercept of the graph of 2x + 3y = 18? slope = -2; y-intercept = (0, 18) slope = negative 2 over 3; y-intercept = (0, 6) slope = negative 3 over 2; y-intercept = (0, 2) slope = 6; y-intercept = (0, negative 2 over 3)

OpenStudy (anonymous):

rewrite the equation in form y=mx+c m is teh slope and c is the y-intercept 2x+3y=18 3y=-2x+18 y=(-2/3)x + 18/3 y = (-2/3)x + 6 So, slope is -2/3 and y-intercept is 6